I need to solve system of equations by addition method
10x+3y=-12
5x=-4y-16
I am lost on this problem also.

Answer by Alan3354(69443)

10x+3y=-12
5x=-4y-16
---------------------
10x+3y=-12
5x+4y =-16
-----------
Multiply the 2nd eqn by -2, then add.